OpenStudy (anonymous):

Hello Mathematicians! Does anyone knows how to solve it? \sin \frac{23\pi}{12}

OpenStudy (jango_in_dtown):

couldnot read it..

OpenStudy (jango_in_dtown):

@Gintare kindly type it once again in the comment section

OpenStudy (anonymous):

\[\sin \frac{ 23\pi }{ 12}\]

OpenStudy (anonymous):

\[\frac{23 \pi }{12}=\frac{23 (180 {}^{\circ})}{12}=345 {}^{\circ} \]

OpenStudy (anonymous):

Yes, but I need to make reduction. So I get sin(270+75), then -cos(30+45) cos(a+b)=cosa*cosb-sina*sinb and the problem is that I cannot make the calculation rigth...

OpenStudy (anonymous):

Writing \(\dfrac{23\pi}{12}=\dfrac{(22+1)\pi}{12}\), you have\[\begin{align*}\sin\frac{23\pi}{12}&=\sin\frac{11\pi}{6}\cos\frac{\pi}{12}+\sin\frac{\pi}{12}\cos\frac{11\pi}{6}\\[1ex] &=-\sin\frac{\pi}{6}\cos\frac{\pi}{12}+\sin\frac{\pi}{12}\cos\frac{\pi}{6}\end{align*}\]Write \(\dfrac{\pi}{12}=\dfrac{\pi}{3}-\dfrac{\pi}{4}\), then \[\begin{align*}\color{white}{\sin\frac{23\pi}{12}}&=-\sin\frac{\pi}{6}\cos\left(\frac{\pi}{3}-\frac{\pi}{4}\right)+\sin\left(\frac{\pi}{3}-\frac{\pi}{4}\right)\cos\frac{\pi}{6}\\[2ex] &=-\sin\frac{\pi}{6}\left(\cos\frac{\pi}{3}\cos\frac{\pi}{4}+\sin\frac{\pi}{3}\sin\frac{\pi}{4}\right)\\[1ex] &\quad\quad+\cos\frac{\pi}{6}\left(\sin\frac{\pi}{3}\cos\frac{\pi}{4}-\sin\frac{\pi}{4}\cos\frac{\pi}{3}\right)\end{align*}\]

OpenStudy (phi):

It's easier (for me) to use degrees so sin(345) we can use -15 instead of 345 sin(345)= sin(-15) and -15 = 30-45 so sin(30-45) = sin(30) cos(45) - cos(30) sin(45) \[ \sin(345)= \frac{1}{2} \frac{\sqrt{2}}{2} - \frac{\sqrt{3}}{2} \frac{\sqrt{2}}{2} \\ = \frac{\sqrt{2}}{4} \left(1 - \sqrt{3}\right) \]

OpenStudy (anonymous):

Thank you very much!!!!